# MENTOR: Mixture-of-Experts Network with Task-Oriented Perturbation for Visual Reinforcement Learning

### v1

@@ Abstract
Visual deep reinforcement learning (RL) enables robots to acquire skills from visual input for unstructured tasks. However, current algorithms suffer from low sample efficiency, limiting their practical applicability. In this work, we present MENTOR, a method that improves both the architecture and optimization of RL agents. Specifically, MENTOR replaces the standard multi-layer perceptron (MLP) with a mixture-of-experts (MoE) backbone, enhancing the agent’s ability to handle complex tasks by leveraging modular expert learning to avoid gradient conflicts. Furthermore, MENTOR introduces a task-oriented perturbation mechanism, which heuristically samples perturbation candidates containing task-relevant information, leading to more targeted and effective optimization. MENTOR outperforms state-of-the-art methods across three simulation domains—DeepMind Control Suite, Meta-World, and Adroit. Additionally, MENTOR achieves an average of 83% success rate on three challenging real-world robotic manipulation tasks including Peg Insertion, Cable Routing, and Tabletop Golf, which significantly surpasses the success rate of 32% from the current strongest model-free visual RL algorithm. These results underscore the importance of sample efficiency in advancing visual RL for real-world robotics. Experimental videos are available at mentor.
Figure 1: MENTOR is validated in real-world tasks. We design three challenging robotic learning tasks for the agent to acquire skills through real-world visual reinforcement learning. MENTOR achieves the most efficient and robust policies compared to the baselines.
@@ 1 Introduction
Visual deep reinforcement learning (RL) focuses on agents that perceive their environment through high-dimensional image data, closely aligning with robot control scenarios where vision is the primary modality. Despite substantial progress in this field (Kostrikov et al., 2020; Yarats et al., 2021; Schwarzer et al., 2020; Stooke et al., 2021; Laskin et al., 2020a), these methods still suffer from low sample efficiency. As a result, most visual RL pipelines have to be first trained in the simulator and then deployed to the real world, inevitably leading to the problem of sim-to-real gap (Zhao et al., 2020; Salvato et al., 2021).
@@
In terms of optimization, visual RL agents often struggle with local minima due to the unstructured nature of robotic tasks. Recent works have shown that periodically perturbing the agent’s weights with random noise can help escape local minima (Nikishin et al., 2022; Sokar et al., 2023; Xu et al., 2023; Ji et al., 2024). However, the choice of perturbation candidates (i.e., the network weights used to perturb the current agent’s weights) has not been thoroughly explored. Building on this idea, we propose a task-oriented perturbation mechanism. Instead of sampling from a fixed distribution, we maintain a heuristically shifted distribution based on the top-performing agents from the RL history. We sample from this heuristic distribution to perturb the current agent’s weights during training. The intuition is that the distribution gradually formed by the weights of previous top-performing agents may accumulate task-relevant information, leading to more promising optimization directions than purely random noise.
@@ 2 Preliminary
The concept of mixture-of-experts (MoE) was first introduced by Jacobs et al. (1991) and Jordan & Jacobs (1994), proposing a simple yet powerful framework where different parts of a model, called experts, specialize in different tasks or different aspects of a task. A sparse MoE layer consists of multiple experts and a router. The router predicts a probability distribution over the experts for a given input. Based on this distribution, only top-$k$ experts are activated for processing the input (Shazeer et al., 2017). Assuming there are $N$ experts, each of which is a feed-forward network (FFN), the final output of the MoE can be written as
@@
where $w(i;\mathbf{x})$ is the gating function determining the utilization of the $i$-th expert for input $\mathbf{x}$. $h(\mathbf{x})$ is the router’s output, producing logits for expert selection, and $\operatorname{topk}\quantity(h(\mathbf{x}))$ selects the top $k$ experts. $\operatorname{softmax}\quantity(\cdot)$ normalizes these top-$k$ values into probabilities.
@@
We employ visual reinforcement learning (RL) to train policies for robotic systems, modeled as a Partially Observable Markov Decision Process (POMDP) defined by the tuple $(S,O,A,P,r,\gamma)$. Here, $S$ is the true state space, $O$ represents visual observations (a stack of three image frames), $A$ is the robot’s action space, $P:S\times A\rightarrow S$ defines the transition dynamics, $r(s,a):S\times A\rightarrow\mathbb{R}$ specifies the reward, and $\gamma\in(0,1]$ is the discount factor. The goal is to learn an optimal policy $\pi_{\theta}(a_{t}\mid o_{t})$ that maximizes the expected cumulative reward $E_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\right]$.
@@
$s_{i}^{l}=\frac{\mathbb{E}_{\bm{x}\in{\mathcal{I}}}|\mathrm{linear}_{i}^{l}(\bm{x})|}{\frac{1}{N^{l}}\sum_{k\in l}\mathbb{E}_{\bm{x}\in{\mathcal{I}}}|\mathrm{linear}_{k}^{l}(\bm{x})|}$ (3)
A neuron $i$ in layer $l$ is considered $\tau$-dormant if its score satisfies $s_{i}^{l}\leq\tau$.
Definition 2.
@@
$\beta_{\tau}=\frac{\sum_{l\in\theta}D_{\tau}^{l}}{\sum_{l\in\theta}N^{l}}$ (4)
As shown by Xu et al. (2023); Ji et al. (2024), the dormant ratio is a critical indicator of neural network behavior and can be leveraged in RL algorithms as an effective metric to improve learning efficiency through parameter perturbation. This process periodically resets the network weights by softly interpolating between the current parameters and randomly initialized values (Ash & Adams, 2020; D’Oro et al., 2022):
$\theta_{k}=\alpha\theta_{k-1}+(1-\alpha)\phi,\quad\phi\sim\text{initializer}$ (5)
Here, $\alpha$ is the perturbation factor, $\theta_{k-1}$ and $\theta_{k}$ are the network weights before and after the reset, respectively, and $\phi$ represents randomly initialized weights (typically drawn from Gaussian noise). The value of $\alpha$ is dynamically adjusted based on the dormant ratio $\beta$ as $\alpha=\text{clip}(1-\mu\beta,\alpha_{\text{min}},\alpha_{\text{max}})$, where $\mu$ is the hyperparameter called the perturbation rate.
@@ 3 Method
3.1 Architecture: Mixture-of-Experts as the Policy Backbone
@@
In challenging robotic learning tasks, RL agents are often assigned $K\geq 2$ different tasks or subgoals, each associated with a loss function $L_{i}(\theta)$. The goal is to find optimal agent weights $\theta\in\mathbb{R}^{m}$ that minimize losses across all objectives. In practice, a common approach is to reduce the average loss over all tasks: $\theta^{*}=\arg\min_{\theta\in\mathbb{R}^{m}}\left\{L_{0}(\theta)\overset{\Delta}{=}\frac{1}{K}\sum_{i=1}^{K}L_{i}(\theta)\right\}$. If the agent uses a shared set of parameters $\theta$ (e.g., MLP), meaning all parameters must be simultaneously active to function, the optimization process using gradient descent may compromise individual loss optimization. This issue, known as conflicting gradients (Yu et al., 2020a; Liu et al., 2021), hinders the agent’s ability to optimize its behavior when facing complex scenarios effectively.
To address this issue, we propose to utilize an MoE architecture as a substitute for the MLP backbone in RL agents. The MoE structure is characterized by its composition of modular experts, $\theta_{\mathrm{MoE}}=\{\theta_{1},\theta_{2},\dots,\theta_{N}\}$, which allows the agent to activate different experts via a dynamic routing mechanism flexibly. This enables gradients from different tasks or subgoals to correspond to different sets of parameters. Specifically, the parameters of a given expert are updated only by gradients from similar task scenarios, thereby effectively alleviating the gradient conflict problem.
As illustrated in Figure 2, the MoE agent first processes visual inputs using a CNN-based encoder, transforming them into a latent space $Z$. The router $h$ computes a probability distribution $h(i\mid z)$ over the experts $i$ for a given latent vector $z\in Z$. The top-$k$ experts are selected based on this distribution, and their softmax weights $w_{i}$ are computed. The outputs $a_{i}$ from these top-$k$ experts are weighted combined to produce the final output $a$, as shown in Equations 1 and 2. This MoE structure enables the agent to route input visual features to specialized experts based on specific objectives, optimizing its performance in challenging scenarios such as multi-tasking or multi-stage processes.
@@
To better illustrate the important role of dynamic modular expert learning for RL agents, we conduct a multi-task experiment (MT5) in Meta-World (MW) (Yu et al., 2020b), where the agent (#$\texttt{\#}\text{Experts}=16,k=4$) is trained to acquire five opposing skills: Open tasks (Door-Open, Drawer-Open, Window-Open) and Close tasks (Drawer-Close, Window-Close). As shown in Figure 3(a), in addition to sharing some experts for handling common knowledge, the Open and Close tasks have their own dedicated experts (Experts 3, 7 for Open and Experts 9, 10 for Close). We evaluate the cosine similarities (Yu et al., 2020a) for both MLP and MoE agents, as shown in Figure 3(b). The MLP’s gradients show significant conflicts between opposing tasks, resulting in a performance gap (100% success for Close tasks, 82% for Open tasks). In contrast, the MoE model demonstrates higher gradient compatibility, achieving 100% success in both task types.
@@
Figure 3: MoE in multi-task scenarios. Left: Expert usage intensity distribution of the MoE agent in opposing tasks. Right: Gradient conflict among opposing tasks for both MLP and MoE agents. The MLP agent frequently encounters gradient conflicts (indicated by negative cosine similarity) when learning multiple skills, while the MoE agent avoids these conflicts (indicated by positive values).
@@
This structural advantage can also be propagated to challenging single tasks, as the dynamic routing mechanism automatically activates different experts to adjust the agent’s behavior throughout the task, alleviating the burden on shared parameters. We illustrate this through training a same-structure MoE agent on a single, highly challenging Assembly task from MW. Figure 4 shows the engagement of the $k=4$ most active experts during task execution, with Expert 15 serving as the shared module throughout the entire policy execution. The other experts vary and automatically divide the task into four distinct stages: Expert 9 handles gripper control for grasping and releasing; Expert 13 manages arm movement while maneuvering the ring; and Expert 14 oversees the assembly process as the ring approaches its fitting location.
@@
Figure 4: MoE in multi-stage scenarios. We present the expert usage intensity during the Assembly task in MW. While Expert 15 remains highly active throughout the entire process, other experts are activated with varying intensity over time, automatically dividing the task into four distinct stages.
3.2 Optimization: Task-oriented Perturbation Mechanism
Neural network perturbation is employed to enhance the exploration capabilities in RL. Two key factors influence the effectiveness of this process $\theta_{k}=\alpha\theta_{k-1}+(1-\alpha)\phi,\phi\sim\Phi$. $\alpha$ is the perturbation factor controlling the mix between current agent and perturbation candidate weights. $\phi$ represents the perturbation candidate sampled from a distribution $\Phi$, which typically is a fixed Gaussian noise $\mathcal{N}(\mu,\sigma)$. Previous works (Sokar et al., 2023; Xu et al., 2023; Ji et al., 2024) have investigated the use of the dormant ratio to determine $\alpha$, resulting in improved exploration efficiency (see Section 2). However, the selection of perturbation candidates has not been thoroughly examined. In this work, we propose sampling $\phi$ from a heuristically updated distribution $\Phi_{\text{oriented}}$, generated from past high-performing agents, to provide more task-oriented candidates that better facilitate optimization.
We define $\Phi_{\text{oriented}}$ as a distribution from which the weights of high-performing agents can be sampled. To obtain this distribution, we dynamically maintain a fixed-size set $S_{\text{top}}=\{(\theta,R)\}$, where $(\theta,R)$ denotes an agent with the weights $\theta$ and achieves episode reward $R$. The desired distribution is approximated using a Gaussian distribution, specifically $\Phi_{\text{oriented}}=\mathcal{N}(\mu_{\theta}^{\text{top}},\sigma_{\theta}^{\text{top}})$, where $\mu_{\theta}^{\text{top}}$ and $\sigma_{\theta}^{\text{top}}$ represent the mean and standard deviation of the weights in $S_{\text{top}}$. As shown in Figure 2, the set $S_{\text{top}}$ is updated during training: at episode $t$, if the agent with weights $\theta_{t}$ achieves reward $R_{t}$ higher than the lowest in $S_{\text{top}}$, the tuple $(\theta_{t},R_{t})$ replaces the one with the lowest reward. This update ensures that $\Phi_{\text{oriented}}$ accurately reflects the current set of high-performing agents, thus providing improved perturbation candidates $\phi$ for future iterations. The pseudocode is shown in Algorithm 1.
@@ 3.2 Optimization: Task-oriented Perturbation Mechanism
1:Initialize the top-performing set $S_{\mathrm{top}}=\emptyset$, perturb interval $\text{T}_{p}$
2:for each episode $t=1,2,\dots$ do
3:     Execute policy $\pi_{\theta_{t}}$ and obtain episode reward $R_{t}$ and dormant ratio $\beta$
4:     if $|S_{\text{top}}|<N$ then
5:         Add $(\theta_{t},R_{t})$ to $S_{\text{top}}$
6:     else if $R_{t}>\min\{R_{i}\ |\ (\theta_{i},R_{i})\in S_{\text{top}}\}$ then
7:         Replace $(\theta_{j},R_{j})=\arg\min\{R_{i}\ |\ (\theta_{i},R_{i})\in S_{\text{top}}\}$ with $(\theta_{t},R_{t})$
8:     end if
9:     if (Number of steps since last perturb) $\geq\text{T}_{p}$ then
10:         Compute mean $\mu_{\theta}^{\text{top}}$ and standard deviation $\sigma_{\theta}^{\text{top}}$ from $S_{\text{top}}$
11:         Sample perturbation candidate $\phi\sim\Phi_{\text{oriented}}=\mathcal{N}(\mu_{\theta}^{\text{top}},\sigma_{\theta}^{\text{top}})$
12:         Calculate perturb factor as in Sokar et al. (2023): $\alpha=\mathrm{clip}(1-\mu\beta,\alpha_{\min},\alpha_{\max})$
13:         Update agent weights: $\theta_{t}=\alpha\theta_{t}+(1-\alpha)\phi$
14:     end if
15:end for

### v3

@@ Abstract
Visual deep reinforcement learning (RL) enables robots to acquire skills from visual input for unstructured tasks. However, current algorithms suffer from low sample efficiency, limiting their practical applicability. In this work, we present MENTOR, a method that improves both the architecture and optimization of RL agents. Specifically, MENTOR replaces the standard multi-layer perceptron (MLP) with a mixture-of-experts (MoE) backbone and introduces a task-oriented perturbation mechanism. MENTOR outperforms state-of-the-art methods across three simulation benchmarks and achieves an average of 83% success rate on three challenging real-world robotic manipulation tasks, significantly surpassing the 32% success rate of the strongest existing model-free visual RL algorithm. These results underscore the importance of sample efficiency in advancing visual RL for real-world robotics. Experimental videos are available at mentor-vrl.
Figure 1: MENTOR validation on real-world tasks. We design three challenging robotic learning tasks where agents are trained from scratch in the real world. MENTOR demonstrates the most efficient and robust policies compared to baseline methods.
Machine Learning, ICML
@@ 1 Introduction
Visual deep reinforcement learning (RL) focuses on agents that perceive their environment through high-dimensional image data, closely aligning with robot control scenarios where vision is the primary modality. Despite substantial progress in this field (Kostrikov et al., 2020; Yarats et al., 2021; Schwarzer et al., 2020; Stooke et al., 2021; Laskin et al., 2020b), these methods still suffer from low sample efficiency. As a result, most visual RL pipelines have to be first trained in the simulator and then deployed to the real world, inevitably leading to the problem of sim-to-real gap (Zhao et al., 2020; Salvato et al., 2021).
@@
In terms of optimization, visual RL agents often struggle with local minima due to the unstructured nature of robotic tasks. Recent works have shown that periodically perturbing the agent’s weights with random noise can help escape local minima (Nikishin et al., 2022; Sokar et al., 2023; Xu et al., 2023; Ji et al., 2024). However, the choice of perturbation candidates (i.e., the network weights used to perturb the current agent’s weights) has not been thoroughly explored. Building on this idea, we propose a task-oriented perturbation mechanism. Instead of sampling from a fixed distribution, we maintain a heuristically shifted distribution based on the top-performing agents from the RL history. The intuition is that the distribution gradually formed by the weights of previous top-performing agents may accumulate task-relevant information, leading to more promising optimization directions than purely random noise.
@@ 2 Preliminary
Mixture-of-experts (MoE), introduced by Jacobs et al. (1991) and Jordan & Jacobs (1994), is a framework where specialized model components, called experts, handle different tasks or aspects of a task. A sparse MoE layer comprises multiple experts and a router. The router predicts a probability distribution over the experts, activating only the top-$k$ for each input (Shazeer et al., 2017). With $N$ experts, each being a feed-forward network (FFN), the MoE output is:
@@
where $w(i;\mathbf{x})$ is the gating function determining the weight of the $i$-th expert for input $\mathbf{x}$, $h(\mathbf{x})$ provides logits for expert selection, and $\operatorname{topk}\quantity(h(\mathbf{x}))$ selects the top $k$.
@@
We employ visual reinforcement learning (RL) to train policies for robotic systems, modeled as a Partially Observable Markov Decision Process (POMDP) defined by the tuple $(S,O,A,P,r,\gamma)$. Here, $S$ is the true state space, $O$ represents visual observations, $A$ is the robot’s action space, $P:S\times A\rightarrow S$ defines the transition dynamics, $r(s,a):S\times A\rightarrow\mathbb{R}$ specifies the reward, and $\gamma\in(0,1]$ is the discount factor. The goal is to learn an optimal policy $\pi_{\theta}(a_{t}\mid o_{t})$ that maximizes the expected cumulative reward $E_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\right]$.
@@
$s_{i}^{l}=\frac{\mathbb{E}_{\bm{x}\in{\mathcal{I}}}|\mathrm{linear}_{i}^{l}(\bm{x})|}{\frac{1}{N^{l}}\sum_{k\in l}\mathbb{E}_{\bm{x}\in{\mathcal{I}}}|\mathrm{linear}_{k}^{l}(\bm{x})|}\ .$ (3)
This neuron is considered $\tau$-dormant if its score $s_{i}^{l}\leq\tau$.
Definition 2.2.
@@
$\beta_{\tau}=\frac{\sum_{l\in\theta}D_{\tau}^{l}}{\sum_{l\in\theta}N^{l}}\ .$ (4)
As shown by Xu et al. (2023); Ji et al. (2024), the dormant ratio is a key metric in neural network behavior and enhances RL efficiency via parameter perturbation. This periodically resets network weights by interpolating between current parameters and random initialization (Ash & Adams, 2020; D’Oro et al., 2022):
$\theta_{k}=\alpha\theta_{k-1}+(1-\alpha)\phi,\quad\phi\sim\text{initializer}\ .$ (5)
Here, $\alpha$ is the perturbation factor, $\theta_{k-1}$ and $\theta_{k}$ are the weights pre- and post-reset, and $\phi$ denotes randomly initialized weights (e.g., Gaussian noise). The value of $\alpha$ dynamically adjusts as $\alpha=\text{clip}(1-\mu\beta,\alpha_{\text{min}},\alpha_{\text{max}})$, where $\mu$ is the perturbation rate and $\beta$ the dormant ratio.
@@ 3 Method
3.1 Mixture-of-Experts as the Policy Backbone
In challenging robotic tasks, RL agents often assigned $K\geq 2$ different tasks or subgoals, each with a loss $L_{i}(\theta)$. The goal is to optimize shared weights $\theta\in\mathbb{R}^{m}$ by minimizing the average loss:
$\theta^{*}=\arg\min_{\theta\in\mathbb{R}^{m}}\left\{L_{0}(\theta)\overset{\Delta}{=}\frac{1}{K}\sum_{i=1}^{K}L_{i}(\theta)\right\}.$
@@ 3.1 Mixture-of-Experts as the Policy Backbone
When using shared parameters $\theta$ (e.g., MLP), meaning all parameters must be simultaneously active to function, the optimization process using gradient descent may compromise individual loss optimization. This issue, known as conflicting gradients (Yu et al., 2020a; Liu et al., 2021), hinders the agent’s ability to optimize its behavior when facing complex scenarios effectively.
We propose replacing the MLP with an MoE backbone. The MoE, composed of modular experts $\theta_{\mathrm{MoE}}=\{\theta_{1},\theta_{2},\dots,\theta_{N}\}$, which allows the agent to activate different experts via a dynamic routing mechanism flexibly. This enables gradients dynamically route from different tasks to specific experts, reducing gradient conflicts. Each expert is updated using gradients from related tasks, addressing conflicts effectively. As shown in Figure 2, the MoE agent uses a CNN encoder to map visual inputs to latent space $Z$. The router $h$ computes a probability $h(i\mid z)$ over experts for latent $z\in Z$. The top-$k$ experts are selected, and their outputs $a_{i}$ are combined using softmax weights $w_{i}$ (Equations 1 and 2). This structure routes inputs to specialized experts, improving multi-task performance.
@@
To illustrate the important role of dynamic modular expert learning for RL agents, we conduct a multi-task experiment (MT5) in Meta-World, training an agent (#$\texttt{\#}\text{Experts}=16,k=4$) for five opposing tasks: Open (Door-Open, Drawer-Open, Window-Open) and Close (Drawer-Close, Window-Close). Figure 3(a) shows Open and Close tasks share some experts but also utilize dedicated ones. To quantitatively demonstrate how much the MoE alleviates the gradient conflict issue, we evaluate the cosine similarities (Yu et al., 2020a) for both MLP and MoE agents in Figure 3(b). The MLP’s gradients show significant conflicts between opposing tasks, while the MoE model demonstrates higher gradient compatibility. As a result, there is a performance gap, with the MLP achieving 100% success in Close tasks but only 82% in Open tasks, whereas the MoE achieves 100% success in both task types.
@@
Figure 3: MoE in multi-task scenarios. Left: Expert usage intensity distribution of the MoE agent in opposing tasks. Right: Gradient conflict among opposing tasks for both MLP and MoE agents. The MLP agent frequently encounters gradient conflicts (indicated by negative cosine similarity) when learning multiple skills, while the MoE agent avoids these conflicts (indicated by positive values). We also provide a comparison of gradient conflicts for MLP and MoE agents in single-task settings, as detailed in Appendix F.
@@
This structural advantage can also be propagated to challenging single tasks, as the dynamic routing mechanism automatically activates different experts to adjust the agent’s behavior throughout the task, alleviating the burden on shared parameters. We illustrate this through training a same-structure MoE agent on a single, highly challenging Assembly task from Meta-World (MW). Figure 4 shows the engagement of the $k=4$ most active experts during task execution, with Expert 15 serving as the shared module throughout the entire policy execution. The other experts vary and automatically divide the task into four distinct stages: Expert 9 handles gripper control for grasping and releasing; Expert 13 manages arm movement while maneuvering the ring; and Expert 14 oversees the assembly process as the ring approaches its fitting location. More detailed results about how MoE alleviates gradient conflicts in the single task are shown in Appendix F.
@@
Figure 4: MoE in multi-stage scenarios. We present the expert usage intensity during the Assembly task in Meta-World. While Expert 15 remains highly active throughout the entire process, other experts are activated with varying intensity over time, automatically dividing the task into four distinct stages.
@@
  Initialize the set $S_{\mathrm{top}}=\emptyset$, perturb interval $T_{p}$
  for each episode $t=1,2,\dots$ do
    Execute policy $\pi_{\theta_{t}}$ and obtain episode reward $R_{t}$ and dormant ratio $\beta$
    if $|S_{\text{top}}|<N$ then
       Add $(\theta_{t},R_{t})$ to $S_{\text{top}}$
    else
       if $R_{t}>\min\{R_{i}\ |\ (\theta_{i},R_{i})\in S_{\text{top}}\}$ then
          Replace $(\theta_{j},R_{j})$ with $(\theta_{t},R_{t})$, where $j=\mathrm{argmin}R_{j}$
       end if
    end if
    if (Number of steps since last perturb) $\geq T_{p}$ then
       Compute mean $\mu_{\theta}^{\text{top}}$ and standard deviation $\sigma_{\theta}^{\text{top}}$ from $S_{\text{top}}$
       Sample perturbation weight $\phi\sim\Phi_{\text{oriented}}=\mathcal{N}(\mu_{\theta}^{\text{top}},\sigma_{\theta}^{\text{top}})$
       Calculate perturb factor as in Sokar et al. (2023): $\alpha=\mathrm{clip}(1-\mu\beta,\alpha_{\min},\alpha_{\max})$
       Update agent weights: $\theta_{t}=\alpha\theta_{t}+(1-\alpha)\phi$
    end if
  end for
3.2 Task-oriented Perturbation Mechanism
Neural network perturbation is employed to enhance the exploration capabilities in RL. Two key factors influence the effectiveness of this process $\theta_{k}=\alpha\theta_{k-1}+(1-\alpha)\phi,\phi\sim\Phi$. $\alpha$ is the perturbation factor controlling the mix between current agent and perturbation candidate weights. $\phi$ represents the perturbation candidate sampled from a distribution $\Phi$, which typically is a fixed Gaussian noise $\mathcal{N}(\mu,\sigma)$. Previous works (Xu et al., 2023; Ji et al., 2024) have investigated the use of the dormant ratio to determine $\alpha$, resulting in improved exploration efficiency (see Section 2). However, the selection of perturbation candidates has not been thoroughly examined. In this work, we propose sampling $\phi$ from a heuristically updated distribution $\Phi_{\text{oriented}}$, generated from past high-performing agents, to provide more task-oriented candidates that better facilitate optimization.
We define $\Phi_{\text{oriented}}$ as a distribution from which high-performing agent weights can be sampled. This distribution is obtained by maintaining a fixed-size set $S_{\text{top}}=\{(\theta,R)\}$, where $(\theta,R)$ represents an agent with weights $\theta$ and episode reward $R$. The distribution is approximated as $\Phi_{\text{oriented}}=\mathcal{N}(\mu_{\theta}^{\text{top}},\sigma_{\theta}^{\text{top}})$, where $\mu_{\theta}^{\text{top}}$ and $\sigma_{\theta}^{\text{top}}$ are the mean and standard deviation of weights in $S_{\text{top}}$. As shown in Figure 2, $S_{\text{top}}$ is updated during training: at episode $t$, if an agent with weights $\theta_{t}$ achieves reward $R_{t}$ exceeding the lowest in $S_{\text{top}}$, $(\theta_{t},R_{t})$ replaces the lowest-reward tuple. This ensures $\Phi_{\text{oriented}}$ reflects current high-performing agents, improving perturbation candidates $\phi$ for subsequent iterations. The pseudocode is in Algorithm 1.